A solid sphere of diameter D = 26 cm has a charge of Q = 4 nano-coulombs uniformly distributed on it. Calculate the magnitude of the electric field at a distance R = 2.6 m from the surface of the sphere
Answer #1

Concept - use the equation for electric field outside a uniformly charged sphere as shown below,3/188.+) = 29.7+ý ser kon!ece) se se she love for uniformly charged sphere, des

***********************************************************************************************
Check the answer and let me know immediately if you find something wrong or missing... I will rectify the mistakes asap if any

for uniformly des charged E= 10 sphere, , for r>R = kanc) (2.6m2 = 15:33 N/C

for uniformly des charged E= 10 sphere, , for r>R = kanc) (2.6m2 = 15:33 N/C